OpenStudy (anonymous):

The area of a rectangular piece of land is 240 square meters. If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points) @jdoe0001

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@kropot72

OpenStudy (jdoe0001):

 \(\bf wl={\color{brown}{ 240}}\implies l=\cfrac{240}{w} \\ \quad \\ (w+3)(l-5)={\color{brown}{ 240}}\implies (w+3)\left(\cfrac{240}{w}-5\right)=240\)